OpenStudy (anonymous):

Expand And Simplify 4 (sigma) n^2/2 n=0 I understand how to do this, but does the N act as a variable starting from one, and expanding to 4?

OpenStudy (anonymous):

replace n by 0, then 1, then 2, then 3, then 4

OpenStudy (anonymous):

oh and put plus signs between the terms

OpenStudy (anonymous):

Okay, thank you :).

OpenStudy (anonymous):

\[\sum_{k=0}^n=a_0+a_1+a_2+...+a_n\] \[\sum_{n=0}^4\frac{n^2}{2}=\frac{0^2}{2}=\frac{1^2}{2}+\frac{2^2}{2}+...\]

OpenStudy (anonymous):

typo there \[\sum_{n=0}^4\frac{n^2}{2}=\frac{0^2}{2}+\frac{1^2}{2}+\frac{2^2}{2}+...\]